Hobbes, the stuffed tiger, has a mass of 28.4 kg. Calvin, the little boy, has a mass of 17.7 kg. In a game of football, Hobbes runs to the right (towards Calvin) at 9.2 m/s. Calvin runs to the left (towards Hobbes) at 12.0 m/s. The two collide and stick together.



What is their final velocity? (Assume that velocities to the right are positive.)
1
Expert's answer
2020-04-16T08:53:50-0400

"m_1=28.4 \ kg, \ v_1=9.2\ m/s"

"m_2=17.7 \ kg, \ v_2=-12.0\ m/s"

"v" - ?


We will use the fact that:

When two objects collide the total momentum before the collision is equal to the total momentum after the collision.

"p_1=m_1v_1" - momentum of Hobbes

"p_2=m_2v_2" - momentum of Calvin

"p_1+p_2" - total momentum (before collision)

"p=mv=(m_1+ m_2)v" - final momentum (after collision)

"p_1+p_2=p"

"m_1v_1+m_2v_2=(m_1+m_2)v"

We have "\ \ v=\frac{m_1v_1+m_2v_2}{m_1+m_2}"

"v=\frac{28.4\times 9.2-17.7\times 12.0}{28.4+17.7}=1.06 \ m/s"

Answer: 1.06 m/s (to the right)